| ID | 
Date | 
Author | 
Group | 
Subject | 
Text | 
  |  
| 
  349  
 | 
20 Sep 2023, 19:42 | 
Peter Zalden | FXE | Issue | Connerning the phase shifter adjustment 
after undulator gap scan via karabo:
 
 
At 19:30 FXE tested the scan change | 
   |  | 
  348  
 | 
20 Sep 2023, 02:09 | 
Jayanath Koliyadu | SPB | Shift end | X-ray delivery
 
 
 
	7 keV, 2.1 mJ, stable delivery
 
 
 
Optical 
laser delivery | 
   |  | 
  347  
 | 
19 Sep 2023, 18:58 | 
Jan Gruenert | PRC | SA1 status | Concerning SASE1:
 
 
Please note that for this week 
the "remote power limit" was set | 
   |  | 
  346  
 | 
18 Sep 2023, 19:14 | 
Yifeng Jiang | FXE | Issue | 
 
	We found that the x-ray pulse energy 
dropped a lot when we changed the undulator 
from 5.5 keV to 5.45 keV. We called BKR to | 
     |  | 
  345  
 | 
15 Sep 2023, 14:54 | 
Chan Kim | SPB | Issue | Vacuum issue at XTD2 tunnel which might 
be triggered by SA1_XTD2_ATT Arm5 (1.2mm 
thick CVD diamond) | 
     |  | 
  344  
 | 
14 Sep 2023, 23:03 | 
Tokushi Sato | SPB | Shift summary | X-ray delivery
 
 
 
	1.1 mj, 564 kHz, 93 pulses, 
stable delivery
 
 
 
Optical | 
   |  | 
  343  
 | 
14 Sep 2023, 16:49 | 
Thomas Baumann |   | Summary: Night 13th and Day 14th | The late shift on the 13th fixed remaining 
problems with the DLD and the calibration 
pipeline. | 
   |  | 
  342  
 | 
14 Sep 2023, 00:05 | 
Tokushi Sato | SPB | Shift summary | X-ray delivery
 
 
 
	1.1 mj, 564 kHz, 93 pulses, 
stable delivery
 
 
 
Optical | 
   |  | 
  341  
 | 
13 Sep 2023, 02:18 | 
Tokushi Sato | SPB | Shift summary | X-ray delivery
 
 
 
	0.5 mj, improved to 1.1 
mj xray energy in the evening
 
 
 
Optical | 
   |  | 
  340  
 | 
09 Sep 2023, 15:03 | 
Naresh Kujala | PRC | Status | Info from Machine RC (Lars):
 
 
----------------------------------------------------------------------------------------------------------------------------------------
 
 
Author: Froehlich | 
   |  | 
  339  
 | 
09 Sep 2023, 11:36 | 
Naresh Kujala | PRC | Status | - Due to the failure of the cryo cold 
compressor. The machine is down.
 
 
- Next update 9 pm today | 
   |  | 
  338  
 | 
09 Sep 2023, 07:54 | 
Naresh Kujala | PRC | Status | Info from MRC (Lars)
 
 
-----------------------------------------------------------------------------
 
 
The cryo crew reports that the | 
   |  | 
  337  
 | 
09 Sep 2023, 07:00 | 
Erik Brambrink | HED | Shift summary | 
 
	Data taken from 11 to cryo failure
 
	Beam 
conditions stable
 
	Timing scan 
for different energy
 
	Operation | 
   |  | 
  336  
 | 
08 Sep 2023, 06:22 | 
Erik Brambrink | HED | Shift summary | 
 
	Improved the alignment procedure:
 
	
 
		Using 
the pin to make sure of the spatial overlap
 
		SAXS 
signal is used to make sure the pin is well | 
   |  | 
  335  
 | 
07 Sep 2023, 23:46 | 
Andreas Scherz | SCS | Shift summary | Shift summary 07.09.2023
 
 
Achievements:
 
1. found t0   on LSCO x=0.15 | 
   |  | 
  334  
 | 
07 Sep 2023, 07:24 | 
Justine Schlappa, Guiseppe Mercurio | SCS | Shift summary, September 7 after nightshift | Achievements:
 
- Optimized hRIXS settings for OK 
edge at 145 deg scattering angle, resolution | 
   |  | 
  333  
 | 
06 Sep 2023, 23:33 | 
Ben, Andreas | SCS | Shift Summary | Achievements
 
 
 
	Determined temporal overlap 
between the X-ray and laser
 
	measured | 
   |  | 
  332  
 | 
06 Sep 2023, 06:35 | 
Giuseppe Mercurio | SCS | Giuseppe, Justine | achievements:
 
 
- we found optimum hRIXS settings 
for 1020 eV but moving the detector arm | 
   |  | 
  331  
 | 
05 Sep 2023, 22:33 | 
Oliver Humphries | HED | Shift summary | 
 
	Received the beam from day shift, 
aligned to beam stop (low signal due to 4m 
air scatter, only 700uJ upstream)
 
	
 
		ZYLA5 | 
   |  | 
  330  
 | 
29 Aug 2023, 03:14 | 
Wonhyuk Jo | MID | Shift summary | Shift summary for 29th of Aug.
 
 
- Conducted Time-resolved Bragg 
CDI experiment using 3rd harmonic pump laser.  | 
   |   
 |